# A Low Rank Neural Representation of Entropy Solutions

### v1

@@ 4 LRNR approximation of classical solutions
In this section, we will discuss how a LRNR can be used to approximate classical solutions to Eq. 2.1. In this case, the solution is given by the method of characteristics, and is known to have an efficient approximation using transported subspaces [70].
We will recall the transported subspaces, and discuss how the inverse that appears in the composition can be avoided straightforwardly using a trick we call the inverse-bias trick. Starting with a given transported subspace in its original form, we will first show that one can rewrite it in a certain discrete form using this trick. Next, we will show that a transported subspace in this discrete form is equivalent to a 2-layer LRNR of comparable complexity. Finally, we will approximate the classical solution to the scalar conservation law using a 2-layer LRNR.
@@ 4.1 The transported subspaces and the inverse-bias trick
We will review the definition of transported subspaces and introduce an approximation technique which allows one to avoid calculating the left-inverses that typically appear in method of characteristics, e.g. in Eq. 2.35. Then we will derive a useful estimate for transport subspaces when this technique is applied.
Throughout this section and the next, we will use the notation $\Omega_{x}$ and $\widehat{\Omega}_{x}$ to denote arbitrary open intervals in $\mathbb{R}$, however in the subsequent sections they will play roles analogous to the spatial domain in Eq. 2.1 and the extended spatial domain Eq. 2.16 above.
@@
Let us define,
(4.10) $f_{0}(x):=\sum_{i=1}^{N}c_{i}\rho(x-x_{i}),\qquad f_{0}\circ^{+}g_{\varepsilon}(x):=\sum_{i=1}^{N}c_{i}\rho(x-g_{\varepsilon}(x_{i})).$
By the triangle inequality, we have
(4.11) $\lVert{f\circ g^{+}-f_{\varepsilon}\circ^{+}g_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$ $\leq\lVert{f\circ g^{+}-f_{0}\circ g^{+}}\rVert_{L^{1}(\Omega_{x})}$
For the first term,
(4.12) $\lVert{f\circ g^{+}-f_{0}\circ g^{+}}\rVert_{L^{1}(\Omega_{x})}$ $=\int_{\Omega_{x}}\left\lvert f(g^{+}(x))-f_{0}(g^{+}(x))\right\rvert\,\mathrm{d}x$
then for the second term,
(4.13) $\lVert{f_{0}\circ g^{+}-f_{0}\circ^{+}g_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$
Next, we have that
(4.14) $\lVert{f_{0}\circ^{+}g_{\varepsilon}-f_{\varepsilon}\circ^{+}g_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$
(4.15) $\lVert{f-f_{0}}\rVert_{L^{1}(\widehat{\Omega}_{x})}$ $\leq\lVert{f-f_{\varepsilon}}\rVert_{L^{1}(\widehat{\Omega}_{x})}+\lVert{f_{\varepsilon}-f_{0}}\rVert_{L^{1}(\widehat{\Omega}_{x})}$
Putting together these inequalities,
and the result follows upon rearranging the terms.
@@ 4.2 LRNR approximation of transported subspaces
In this section, we will show that a 2-layer version of LRNR, a simple special case of the general definition of LRNRs introduced in Section 3, is able to approximate any transported subspace. The rank of the approximating LRNR is equivalent to the dimension of the transported subspace.
Example 4.5 (A 2-layer version of LRNR).
Suppose we are given
@@
We now show that every transported subspace $\bar{H}_{r-1}$ can be uniformly approximated by a LRNR $H_{r}$ with comparable rank. LRNR needs one more rank due to a technicality in representing the jump function using Eq. 4.6.
@@
Proof 4.7.
(4.23) $f(\beta)=\sum_{i}\beta_{i}\psi_{i}=\sum_{k=1}^{K}\bar{c}_{k}(\beta)\rho(\,\cdot-x_{k}),$
(4.24) $\lVert{f(\beta)\circ g(\alpha)^{+}-f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha)}\rVert_{L^{1}(\Omega_{x})}$
(4.25) $\lVert{f(\beta)\circ g(\alpha)^{+}-f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha)}\rVert_{L^{1}(\Omega_{x})}\leq\left\lvert{f(\beta)}\right\rvert_{TV(\widehat{\Omega}_{x})}\lVert{\rho-\rho_{\varepsilon}}\rVert_{L^{1}(\mathbb{R})}.$
(4.26) $f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha)(x)$
$=\sum_{k=1}^{K}\bar{c}_{k}(\beta)\rho_{\varepsilon}(x-g(\alpha)(x_{k}))$
so one can write
(4.27) $f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha)(x)=W^{2}\sigma(W^{1}x+B^{1})+B^{2},\qquad x\in\Omega_{x},$
with the weights and biases
(4.28) $W^{1}$ $=\left[1,1,\cdots,1\right]^{T}\in\mathbb{R}^{2K\times 1},$
$B^{2}$ $=0\in\mathbb{R}^{1\times 1}.$
We reorganize the NNs as 2-layer LRNR $H_{r}$ in the form Eq. 4.20 in Example 4.5. Let $\{e_{i}\}_{i=1}^{r}$ denote the standard basis of $\mathbb{R}^{r}$. We write the weights $W^{1},W^{2}$ as
(4.29) $W^{1}=e_{1}\cdot\mathbb{U}_{r}^{1},\quad W^{2}=[\beta_{1},\cdots,\beta_{r-1},0]\cdot\mathbb{U}_{r}^{2},$
(4.30) $\mathbb{U}_{r}^{1}$ $=(U_{i}^{1})_{i=1}^{r},$ $\left[U_{i}^{1}\right]_{j}:=1,\text{ for all }i,j,$
Write the biases $B^{1},B^{2}$ as
(4.31) $B^{1}=[\alpha_{1},\alpha_{2},\cdots,\alpha_{r-1},1]\cdot\mathbb{V}^{1},\quad B^{2}=0\cdot\mathbb{V}^{2},$
(4.32) $\mathbb{V}_{r}^{1}$ $=(V_{i}^{1})_{i=1}^{r},$$\begin{cases}\left[V_{i}^{1}\right]_{j}:=\psi_{i}(x_{\lfloor j/2\rfloor})&\text{ for }i=1,...\,,r-1,\\ \left[V_{i}^{1}\right]_{j}:=0&\text{ for }i=r,\end{cases}$
$\mathbb{V}_{r}^{2}$ $=(V_{i}^{2})_{i=1}^{r},$ $V_{i}^{1}:=0,\quad\text{ for all }i.$
(4.33) $\gamma^{1}$ $=e_{1},$ $\gamma^{2}$ $=[\beta_{1},\cdots,\beta_{r-1},0]^{T},$
$\theta^{1}$ $=[\alpha_{1},\cdots,\alpha_{r-1},1]^{T},$ $\theta^{2}$ $=0,$
the LRNR $h(\gamma,\theta)$ is defined as in Example 4.5 with the coefficients Eq. 4.33 and weights and biases Eqs. 4.29 and 4.31. By construction, it satisfies
(4.34) $h(\gamma,\theta)=f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha).$
Plugging Eq. 4.34 into Eq. 4.25,
(4.35) $\lVert{f(\beta)\circ g(\alpha)^{+}-h(\gamma,\theta)}\rVert_{L^{1}(\Omega_{x})}\leq\left\lvert{f(\beta)}\right\rvert_{TV(\widehat{\Omega}_{x})}\lVert{\rho-\rho_{\varepsilon}}\rVert_{L^{1}(\mathbb{R})}.$
The desired inequality follows from
(4.36) $\left\lvert{f(\beta)}\right\rvert_{TV(\widehat{\Omega}_{x})}\leq\left\lvert{\sum_{i}\beta_{i}\psi_{i}}\right\rvert_{TV(\widehat{\Omega}_{x})}\leq\max_{i}\left\lvert\beta_{i}\right\rvert\sum_{i}\left\lvert{\psi_{i}}\right\rvert_{TV(\widehat{\Omega}_{x})}.$
@@ 4.3 LRNR approximation of classical solutions
there is a LRNR $H_{r}$ with rank $r=3$, d.o.f $K$, width $M\sim K$, and depth $L=2$ for every given $\varepsilon>0$ such that, for coefficients $\gamma(t),\theta(t)$ that are affine functions of $t$, the member $h(t)=h(\gamma(t),\theta(t))\in H_{r}$ achieves the uniform error
@@
Proof 4.9.
We write the classical solution $u(t)=u(\cdot,t)$ as
(4.39) $u(\cdot,t)$ $=u_{0}\circ X^{-1}(\cdot,t)$
then it is clear that the set of solutions forms a transported subspace, so Theorem 4.6 applies. Recall that the approximation proceeds as follows: Letting $\widehat{\Omega}_{x}=\Omega_{x}$, define $u_{0\varepsilon},X_{\varepsilon}\in C(\Omega_{x})\cap P_{1}(\Omega_{x})$ for which Lemma 4.3 implies
(4.40) $\lVert{u_{0}\circ X^{+}-u_{0\varepsilon}\circ^{+}X_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$
(4.41) $\lVert{u_{0}\circ X^{+}-u_{0\varepsilon}\circ^{+}X_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$
As shown in Eq. 4.34 in the proof of Theorem 4.6, there is a LRNR $h$ satisfying
(4.42) $\lVert{u_{0\varepsilon}\circ^{+}X_{\varepsilon}(t)-h(t)}\rVert_{L^{1}(\Omega_{x})}$
due to the estimate
(4.43) $\lVert{X^{\prime}}\rVert_{L^{\infty}(\Omega_{x})}\leq 1+T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\lVert{u^{\prime}}\rVert_{L^{\infty}(\Omega_{x})}.$
Setting $\varepsilon$ to be sufficiently small, the result follows by Eq. 4.41 and Eq. 4.42.
The coefficients $\gamma,\theta$ are induced by the affine mapping $\mu$ given by Theorem 4.6, see Eq. 4.33. The transported subspace coefficients depend linearly on $t$, thus $\gamma,\theta$ are affine functions of $t$.
@@ 5 LRNR approximation of entropy solutions
In this section, we construct a LRNR approximation of the entropy solution. We will first provide a broad outline in Section 5, then work towards the main results Theorems 5.5 and 5.7 by preparing the prerequisites in Sections 5.2 and 5.3.
@@ 5.1 Outline of approach
In light of the inverse-bias trick in Section 4.1, it might appear that the approximation of the entropy solution is now a simple matter of approximating the rarefied characteristics $\widehat{X}$ Eq. 2.22. Unfortunately, the rarefied characteristics manifold given by
@@
itself has a slowly decaying Kolmogorov width in the presence of shocks. In plain terms, this is due to the fact that the endpoints of the constant region of $\widehat{X}(\cdot,t)$ describing the shocks move with time, according to the Rankine-Hugoniot shock speed (see Figures 5 and 6, and comments in Section 2.4). This leads to sharp features (kinks) that travel across the domain $\widehat{\Omega}_{x}$ as time evolves [71]. This is an obstruction to the strategy of finding a compositional reduced model directly, by using e.g. transported subspaces (Definition 4.2). Put in other terms, there is now a slow decay in the Kolmogorov width of both the solution manifold and the rarefied characteristics manifold.
One natural question to ask is whether it is possible to adopt the general strategy used to construct compositional reduced models yet again, this time for the manifold of rarefied characteristics $\mathcal{M}_{\widehat{X}}$. A natural compositional structure is given by the shock time function $\lambda$ Eq. 2.29, and a low-rank structure is apparent in the relief characteristics $\widecheck{X}$ Eq. 2.33, in the superpositions of a few basis functions; see the diagram in Eq. 3.1. To provide a succinct analogy: The classical characteristics are to the solution manifold, what the shock time function is to the rarefied characteristics manifold.
We need care in introducing the shock-related jump term $\rho(\lambda(x)-t)$ that appears in the inner-most composition Eq. 2.33. In the relief characteristics, this term is used in conjunction with the left-inverse to replace the portions of the rarefied characteristic curves with constant regions. LRNRs are defined using ReLU activations so they are continuous piecewise linear functions; as such, the jump discontinuity must be approximated.
We will show that it is still possible to construct an analogous approximation. We aim to (1) replace certain parts of the rarefied characteristics $\widehat{X}$ Eq. 2.22 in a manner similar to that of relief characteristics $\widecheck{X}$ Eq. 2.33, (2) without relying on the left-inverse, and (3) without using piecewise constant jump function $\rho$ Eq. 2.32. To this end, we substitue $\rho$ with $\rho_{\varepsilon}$ Eq. 4.6 then devise a way to apply the inverse-bias trick. We start with a continuous piecewise linear approximation of the rarefied characteristics with multivalued inverse Eq. 2.27,
@@
Then, we represent shocks (constant regions) by removing certain variations in the characteristics, emulating the relief characteristics while preserving continuity and piecewise linearity: We introduce a layered approximation so that slices of $\widehat{X}_{0}(\cdot,t)$ are translated outside of the domain. To maintain uniform error between $\widehat{X}$ and this approximation, it is necessary that the slice that is being translated to the right does so quickly and is small in amplitude. A key observation that makes this possible is that, in the approximation Eq. 5.2, there is a freedom in constructing the coefficients $(c_{k})$. Without adding to the number of coefficients, one may restructure the coefficients so that the approximation is rewritten as a sum of the slices $(\eta_{k})$,
(5.3) $\widehat{X}_{0}(\cdot,t)\approx\sum_{k}t\,\eta_{k}\left(\,\cdot\,-x_{k}\right),\qquad\lVert{t\eta_{k}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\lesssim\varepsilon.$
Now each slice $\eta_{k}$ can be moved out of the domain using the inverse-bias trick devised for Eq. 2.33: This amounts to inserting an approximation of the term $\rho(t-\lambda(x))$ into the bias,
(5.4) $\widehat{X}(\cdot,t)\approx\sum_{k}c_{k}\rho_{\varepsilon}\left(\,\cdot\,-C_{x}\rho_{\varepsilon}(t-\lambda(x_{k}))-x_{k}\right),\qquad C_{x}\geq\left\lvert\widehat{\Omega}_{x}\right\rvert+\varepsilon.$
Whenever $|t-\lambda(x_{k})|\lesssim\varepsilon$, i.e. approximately the time when the characteristic curve emanating from $x_{k}$ has entered into the shock, a portion of the characteristic information $\eta_{k}$ is moved outside of the domain $\widehat{\Omega}_{x}$ with speed $\sim 1/\varepsilon$. The approximation maintains the monotonicity of the characteristics only approximately, but it is still well-defined where the approximation is not monotone thanks to the inverse-bias trick (Section 4.1).
We present our construction as follows. In Section 5.2, we derive a time-layered approximation of rarefied characteristics $\widehat{X}$ that is valid for short time intervals, then in Section 5.3 we construct a discretization of the form Eq. 5.4, then finally in Section 5.4 we present the LRNR approximation result.
@@
We close this outline with the remark that the use of a threshold function in Eq. 5.4 shares a similar theme with TSI [82], transport reversal [69], or FTR [43].

### v2

@@ 4 LRNR approximation of classical solutions
In this section, we will discuss how a LRNR can be used to approximate classical solutions to (2.1). In this case, the solution is given by the method of characteristics, and is known to have an efficient approximation using transported subspaces [71]. For the classical case, it is simpler to suspend the assumption that the initial condition is piecewise constant with finitely many jumps (that is, $u_{0}\in\overline{\mathcal{U}}$). Then, for a smooth initial condition $u_{0}$ the classical solution is
(4.1) $u(x,t)=u_{0}\circ X^{-1}(x,t),$
with classical characteristics $X$ that solve the ODE (simpler version of (2.4)),
(4.2) $\frac{\partial X}{\partial t}=F^{\prime}(u(X,t)),\quad X(x,0)=x,\quad(x,t)\in\Omega_{x}\times\Omega_{t}.$
Due to the absence of rarefaction and shock waves in the classical case, the solutions simplifies in two major ways: (1) The inverse $(\cdot)^{-1}$ is used in the solution (4.1) rather than the left-inverse $(\cdot)^{+}$ used in (2.25, 2.33), and (2) the initial condition $u_{0}$ and the characteristic curves $X$ are directly used instead of their extensions (i.e. $\hat{u}_{0}$, $\widehat{X}$ or $\widecheck{X}$).
We focus on the fact that the classical solution (4.1) resembles the compositions of superpositions (3.7) we have encountered when deriving LRNRs in the previous section. We may write the classical solution in a similar form,
(4.3) $u(x,t)$ $=\left(\sum_{i=1}^{r}\gamma^{2}_{i}(t)\phi_{i}\right)\circ\left(\sum_{i=1}^{r}\gamma^{1}_{i}(t)\varphi_{i}\right)^{-1}(x)$
[…]
An important difference with (3.7) is the presence of the inverse $(\cdot)^{-1}$. This must be carefully handled to approximate this compositional form of the classical solution using LRNRs. This issue arose in the compositional reduced models and other neural network approximations [78, 47, 71]. Our discussion here centers specifically on transported subspaces [71] which are most directly related to LRNRs.
We will recall the transported subspaces, and discuss how the inverse that appears in the composition can be avoided straightforwardly using a trick we call the inverse-bias trick. Starting with a given transported subspace in its original form, we will first show that one can rewrite it in a certain discrete form using this trick. Next, we will show that a transported subspace in this discrete form is equivalent to a 2-layer LRNR of comparable complexity. Finally, we will approximate the classical solution to the scalar conservation law using a 2-layer LRNR. The techniques developed in this section for approximating the classical solutions will be used throughout the subsequent Section 5 that addresses the entropy solution. We refer the reader to Appendix A for a summary of notions including the transported subspaces, the inverse-bias trick, and the LRNR approximation of the classical solution introduced in this section.
@@ 4.1 The transported subspaces and the inverse-bias trick
We will review the definition of transported subspaces and introduce an approximation technique which allows one to avoid calculating the left-inverses that typically appear in method of characteristics, e.g. in (2.34). Then we will derive a useful estimate for transport subspaces when this technique is applied.
Throughout this section and the next, we will use the notation $\Omega_{x}$ and $\widehat{\Omega}_{x}$ to denote arbitrary open intervals in $\mathbb{R}$, however in the subsequent sections they will play roles analogous to the spatial domain in (2.1) and the extended spatial domain (2.16) above.
@@
See Appendix B.2.
@@ 4.2 LRNR approximation of transported subspaces
In this section, we will show that a 2-layer version of LRNR, a simple special case of the general definition of LRNRs introduced in (3), is able to approximate any transported subspace. The rank of the approximating LRNR is equivalent to the dimension of the transported subspace.
{expl}
[A 2-layer version of LRNR] Suppose we are given
@@
We now show that every transported subspace $\bar{H}_{r-1}$ can be uniformly approximated by a LRNR $H_{r}$ with comparable rank. LRNR needs one more rank due to a technicality in representing the jump function using (4.9).
@@
Proof 4.6.
See Appendix B.3.
@@ 4.3 LRNR approximation of classical solutions
there is a LRNR $H_{r}$ with rank $r=3$, d.o.f $K$, width $M\sim K$, and depth $L=2$ for every given $\varepsilon>0$ such that, for coefficients $\gamma(t),\theta(t)$ that are linear or constant functions of $t$, the member $h(t)=h(\gamma(t),\theta(t))\in H_{r}$ achieves the uniform error
@@ 5 LRNR approximation of entropy solutions
Having constructed an efficient LRNR approximation of the classical solution, we turn our attention to the LRNR approximation of the entropy solution. As briefly discussed in the beginning of Section 4, there are important differences between the classical solution and the entropy solution. Let us explain how these differences introduce significant challenges in constructing the LRNR approximation.
Recall from Section 2 that we have introduced two new representations of the entropy solution, namely, one using the rarefied characteristics $\widehat{X}$ (2.25) and another using the relief characteristics $\widecheck{X}$ (2.33):
(5.1) $u_{0}\circ X^{-1}(x,t)$ classical characteristics,
$\hat{u}_{0}\circ\widehat{X}^{+}(x,t)$ rarefied characteristics,
$\hat{u}_{0}\circ\widecheck{X}^{+}(x,t)$ relief characteristics.
$\quad=\hat{u}_{0}\circ(\hat{I}+t(F^{\prime}\circ\hat{u}_{0})+C_{x}\rho\circ(t-\lambda))^{+}$
The rarefied characteristics $\widehat{X}$ is a direct generalization of the classical chracteristics that allows representing shocks and rarefaction waves. It is monotone, thus the inverse-bias trick can still be applied directly to handle the left-inverse $(\cdot)^{+}$. However, as we discuss below, the rarefied characteristics are not in a compositional low rank form that is compatible with an efficient representation by LRNRs. On the other hand, the relief chracteristics $\widecheck{X}$ do have a compositional low rank structure (as discussed in Section 3) that is compatible with LRNRs. However, it is not monotone and hence the left-inverse cannot be removed using the inverse-bias trick directly.
Therefore, our goal here is to devise a LRNR approximation that agrees with the rarefied characteristics $\widehat{X}$, while simultaneously exploiting the compositional low-rank structure of the relief characteristics $\widecheck{X}$. We build a construction by decomposing the relief characteristics into sum of local-in-time contributions, and introduce a LRNR structure per each summand via the same technique used to for relief characteristics. This allows us to take advantage of the compositional structure while approximately maintaining monotonicity.
In this section we will address these challenges involved and construct an efficient LRNR approximation of the entropy solution. We first provide a broad outline in Section 5.1, then introduce various discretizations that gradually reveal the neural-network-like structure in Sections 5.2 and 5.3, to finally work towards the main Theorem 5.7 in Section 5.4. Throughout, we refer the reader to Appendix A for a glossary of various approximations introduced in this section.
@@ 5.1 Outline of approach
Figure 7: The time evolution of the rarefied characteristics $\widehat{X}(\cdot,t)$ for the stationary shock example in Figure 5, zoomed into the region relevant to shock propagation. The constant interval (marked by $\Upsilon_{t}$) expands over time $0<t_{1}<t_{2}<t_{3}$. The two kinks at the endpoints (marked by two arrows) of this interval travel in opposite directions over time.
In this section, we will motivate and outline our general approximation approach. We begin by detailing the technical challenges mentioned in the introduction of this section. In light of the inverse-bias trick in Section 4.1, it might appear that the approximation of the entropy solution is now a simple matter of approximating the rarefied characteristics $\widehat{X}$ (2.22). Unfortunately, the rarefied characteristics manifold given by
@@
itself has a slowly decaying Kolmogorov $n$-width in the presence of shocks. For example, recall the stationary shock example from Figure 5. Observe how the rarefied characteristics $\widehat{X}(\cdot,t)$ evolves over time, as illustrated in Figure 7. The constant region of $\widehat{X}(\cdot,t)$ describing the shock expand with time, according to the Rankine-Hugoniot condition, and this leads to kinks that travel across the domain $\widehat{\Omega}_{x}$. As a result, the $n$-width in $L^{2}$-norm must be $\gtrsim n^{-3/2}$ [83, 60, 31, 72]. This is an obstruction to the strategy of finding a compositional reduced model directly, by using e.g. transported subspaces (Definition 4.2). In other words, there is now a slow decay in the $n$-width of both the solution manifold and the rarefied characteristics manifold. [^2] To provide a succinct analogy: The classical characteristics are to the solution manifold, what the shock time function is to the rarefied characteristics manifold.
One natural question to ask is whether it is possible to construct compositional reduced models yet again, this time for the manifold of rarefied characteristics $\mathcal{M}_{\widehat{X}}$. A low-rank structure is apparent in the relief characteristics $\widecheck{X}$ (2.32), in the superpositions of a few basis functions, as discussed in Section 3; see the diagram in (3.1). Still, we need care in introducing the shock-related jump term $\rho(\lambda(x)-t)$ that appears in the inner-most composition (2.32). In the relief characteristics, this term is used in conjunction with the left-inverse to replace the portions of the rarefied characteristic curves with constant regions. LRNRs are defined using ReLU activations so they are continuous piecewise linear functions; as such, the jump discontinuity must be approximated carefully so that the low-rank compositional structure is preserved. We propose a LRNR approximation that satisfies these requirements, outlined as follows.
  • We start with a continuous piecewise linear approximation of the rarefied characteristics with multivalued inverse (2.26),
@@
    This approximation so far does not properly handle shocks.
  • As discussed above, the LRNR approximation of shock propagation must be done carefully to preserve the compositional low rankness of the overall approximation and avoid the slow decay in the Kolmogorov width. We represent shocks (constant regions) by removing certain variations in the characteristics, emulating the relief characteristics while preserving continuous piecewise linearity and monotonicity. Roughly speaking, we achieve this by introducing a time-layered approximation so that slices of $\widehat{X}_{0}(\cdot,t)$ are translated outside of the domain.
    To maintain uniform error between $\widehat{X}$ and this approximation, it is necessary that the slice that is being translated to the right does so quickly and is small in amplitude. A key observation that makes this possible is that, in the approximation (5.3), there is a freedom in constructing the coefficients $(c_{k})$. Without adding to the number of coefficients, one may restructure the coefficients so that the approximation is rewritten as a sum of the slices $(\eta_{k})$. With some simplifications, one writes
    (5.4) $\widehat{X}_{0}(\cdot,t)\approx\hat{I}(\cdot)+tv(\cdot)+\sum_{k}t\,\eta_{k}\left(\,\cdot\,\right),\qquad\lVert{t\eta_{k}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\lesssim\varepsilon.$
    Note that this approximates the multi-valued inverse and does not properly represent shocks.
  • We modify the above approximation to represent shocks, by an appropriate removal of slice terms to mimick the expanding constant regions of $\widehat{X}$. That is, we write
    (5.5) $\widehat{X}(\cdot,t)\approx\barwedge(\cdot,t)\approx\hat{I}(\cdot)+tv(\cdot)+\sum_{\mathclap{\begin{subarray}{c}\text{partial sum}\\ \text{over }k\end{subarray}}}t\,\eta_{k}\left(\,\cdot\,\right),\qquad\lVert{t\eta_{k}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\lesssim\varepsilon.$
@@
  • Now each slice $\eta_{k}$ can be removed from the domain using the inverse-bias trick devised for (2.32): This amounts to inserting an approximation of the term $\rho(t-\lambda(x))$ into the bias,
    (5.6) $\widehat{X}(\cdot,t)\approx\doublebarwedge(\cdot,t)\approx\hat{I}(\cdot)+tv(\cdot)+\sum_{k}t\eta_{k}\left(\,\cdot\,-C_{x}\rho_{\varepsilon}(t-\lambda(x_{k}))\right),\quad x_{k}\in\operatorname{supp}\eta_{k},$
[…]
We present our construction as follows. In Section 5.2, we derive a time-layered approximation of rarefied characteristics $\widehat{X}$ that is valid for short time intervals, then in Section 5.3 we construct a discretization of the form (5.6), then finally in Section 5.4 we present the LRNR approximation result.
[…]
We close this outline with the remark that the use of a threshold function in (5.6) shares a similar theme with TSI [83], transport reversal [70], or FTR [44].